Find the equation of tangents to
x^2+y^2-6x-2y+9=0
through the origion.

Answer by stanbon(75887)

Find the equation of tangents to
x^2+y^2-6x-2y+9=0
through the origin.
-------------------------
Take the derivative to get:
2x+2yy'-6-2y'=0
y'(2y-2) = (6-2x)
y' = (6-2x)/(2y-2)
y'(0,0) = 6/-2 = -3
--------------------------
Equation of the tangent at (0,0) is y = -3x
==========================
Cheers,
Stan H.